If there are 10 people in a room what is the probability that any 2 people have the same birthday?
11.7%. (But when there are 23 people, the odds are 50.7%.) In order for 2 people to have the same birthday, then there can only be 9 different dates out of 365 (actually 37 out of 1461). This is roughly equivalent to 2.53% of possible dates. However, the chance of having matching birthdays to any other birthday is actually much larger, because EACH individual has up to 9 chances out of 365 to match another. The probability is calculated using the product of the series p(n) = 1-(n-1/365), which compares the likelihood of all possible pairs of identical birthdays. The chance that no two share a birthday (including the leap day) can be calculated as the product of the terms n = 1 to 10 of [(1461-(4n+1)/1461]. This should also equate to a likelihood of 88% because the Leap Day is only 1 chance in 1461. (see related link)
What is the probability of Down syndrome if the parents are Asian?
the probobility for downsyndrome depends if any of the parents had it or if a relative had it. it doesn't matter if you are Asian or not. you might get it or you may not. it all depends if you or your relatives have it.
How many ways can 5 people sit in a row if two of them insist on sitting together?
(If they wouldn't want to sit together the answer would be 5! (120) but since there is a pair, we can count them as one person. So, my guess is that the answer is 4! (24))
The person who wrote this answer thought correctly that we need to consider the two person as if they are one so the answer is 4! , but still those two person can switch places between themselves so the correct answer should be 4! * 2! = 48
A number cube is rolled three times. What is the probability of rolling a 2 each time?
the chances of rolling a 2 each time during 1 set of 3 rolls is -1 in 216- sets. i believe the formula is: 1 in 6 - the chance of rolling a"2" then multiply the chances for each separate roll 1/6 x 1/6 x 1/6 = 1/216
The odds of any card pulled from an ordinary deck of 52 cards being an Ace is 4 in 52 (4 aces in a deck of 52). This can be reduced to a 1 in 13 chance of any random card pulled from the deck being an Ace (or any other specific value, for that matter).
That 13th last card dealt in a hand is no different than picking a random card out of the pack, regardless of what cards you deal before (face down or blindfolded or even face up, it doesn't matter).
A more interesting question would be "what would the probability be of ANY of those 13 cards being an Ace?" Any takers?
It is 1/2 * 1/2 = 1/4
The probability of drawing 3 red marbles in a row from a bag of 3 blue and 4 red marbles without replacement is right at about 0.11 or about 11 in 100, if you want odds.
Remember that probability is a pure number between zero (no chance - it's impossible) and 1 (it will happen - it's a certainty). A probability of 0.5 is a 50-50 shot, like a coin toss. We good? Let's move on.
We have 7 marbles, and there is a 4 in 7 shot of getting a red marble in the first draw. There is a 3 in 6 shot the next draw since now we only have 6 marbles total (NO replacement) and only 3 red ones (we already drew a red one out). Similarly, there is a 2 in 5 shot of getting a third red one if the first two are red. The probability of the first action (getting a red marble on the first draw) is 4/7 or a probability of about 0.5714. Second draw, it's 3/6 or a probability of 0.5. Last, it's 2/5 or a probability of 0.40. To find the intersection of these probabilities, we just multiply them all together, and we'll get right at about 0.011 for a probability.
What is negative 2 plus negative 8?
well since they have the same signs you add them together and since 2 plus 8 is ten the answer is -10
What is the probability of getting at least 2 heads in three flips of a fair coin?
In three flips of a fair coin, there are a total of 8 possible outcomes: T, T, T; T, T, H; T, H, T; T, H, H; H, H, H; H, H, T; H, T, H; H, T, T Of the possible outcomes, four of them (half) contain at least two heads, as can be seen by inspection. Note: In flipping a coin, there are two possible outcomes at each flipping event. The number of possible outcomes expands as a function of the number of times the coin is flipped. One flip, two possible outcomes. Two flips, four possible outcomes. Three flips, eight possible outcomes. Four flips, sixteen possible outcomes. It appears that the number of possible outcomes is a power of the number of possible outcomes, which is two. 21 = 2, 22 = 4, 23 = 8, 24 = 16, .... Looks like a pattern developing there. Welcome to this variant of permutations.
What is the probability of pulling a queen out of a stack of 52 cards?
The probabilty of pulling a queen out of a stack of cards is four chances (since there are four queens, one of each suit) out of 52. Or 4/52, or .08, or 8%.
What is the probability of drawing 13 cards of different values from a deck of 52?
It is (413*13!)/(52C13) = 0.0001057 approx.
What is the probability that the player will beat the dealer in blackjack?
If you mean, on a specific RANDOM hand, there is, depending on the precise rules, approx. a 43% chance of winning the hand, a 9% chance of a tie, and and a 48% chance of losing the hand.
What is the probability that the dealer will not get blackjack?
The chance of a Blackjack appearing for any player OR the dealer is slightly different depending solely upon how many decks are used.
The approximately correct answer in all cases though is odds of 20-1
That is to say, that one out of every 21 hands is a Blackjack.
Ergo, the probability of the occurrence is approx. 4.8%
The question asked is "What is the probability of the Dealer not getting a blackjack."
Therefore, the correct answer is 100% - 4.8% = 95.2%
If you consider that there is 1 Ace per suit, and 4 tens per suit, then the correct answer is (1/13 * 4/13) = 4/169 or a 2.367% chance that the dealer will get blackjack.
The chance that with an Ace showing that the dealer has a 10 under the Ace, is 4/13 or 30.769%.
With a 10 showing, the chance the dealer has an Ace below the 10 is 1/13 or 7.692%.
The reason you aren't supposed to take insurance, is because of the 30% chance to have it. However if the deck is +6, then even a professional card counter will take insurance.
What is the least number that has three different prime factors?
If by "least number" you mean "smallest positive integer", then the answer is the product of the three smallest primes: 2x3x5 = 30
What is the possibility of tossing heads on each of first 2 tosses?
It is certainly possible.
The probability is 1/4.
When a coin is tossed 9 times in a row what is the probability that heads comes up 9 times?
It is (1/2)9 = 1/512
Why the normal distribution is important in public health research?
I just want to have the question answered for class
What is the probability of NOT drawing a spade from a regular deck of cards?
On a random draw of a single card, the probability is 3/4.